# Causal Inference in Longitudinal Data under Unknown Interference

### v4

@@ 4.1 Asymptotic Distribution
We restrict the dependence in $Y_{i}$ by assuming approximate neighborhood interference (ANI) introduced in Leung (2022). Let unit $i$’s $p$-neighborhood in $\mathcal{G}_{N}$ be defined as $\mathcal{N}_{\mathcal{G}_{N}}(i,p)$, the set of units whose proximity to $i$ does not exceed $p$. We denote an independent copy of the treatment history outside $i$’s $p$-neighborhood as $\mathbf{A}^{(1:T)}_{i,p}=\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{\mathcal{G}_{N}}(i,p)},\tilde{\mathbf{A}}^{(1:T)}_{\mathcal{N}_{N}\setminus\mathcal{N}_{\mathcal{G}_{N}}(i,p)}\right)$. We then impose the following assumption:
@@
$\max_{i\in\mathcal{N}_{N}}E\Bigg{|}Y_{i}\left(\mathbf{A}^{(1:T)}\right)-Y_{i}\left(\mathbf{A}^{(1:T)}_{i,p}\right)\Bigg{|}\leq\theta_{N,p},$
@@
This assumption implies that for any bounded Lipschitz function $L(\cdot)$, the covariance between $L\left(\mathbf{Z}_{i}(d)Y_{i}\right)$ and $L\left(\mathbf{Z}_{j}(d)Y_{j}\right)$ can be bounded by a transformation of $\theta_{N,p}$. This property, known as $\psi$-dependence (Doukhan and Louhichi, 1999; Kojevnikov, Marmer and Song, 2021), ensures that the dependence between outcomes decays as the proximity between units increases, even though it may never disappear.
@@
$M_{N}(q,k)=\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q)\right|^{k},\vspace{-1em}$
@@
There exists a constant $\tilde{y}<\infty$ such that
$|Y_{i}\left(\mathbf{a}^{(1:T)}\right)|\leq\tilde{y}\vspace{-1em}$
@@
We discuss potential relaxations of this assumption in Section 5. Additionally, we assume that the propensity score can be consistently estimated at a parametric rate (root-$N$ or faster) in observational studies:
@@
[…]
@@
Under Assumptions 1-6, if there exist $\epsilon>0$ and a sequence $\{q_{N}\}$ that grows with $N$ such that
$\frac{N^{\frac{3}{2}}}{\Sigma_{N}^{\frac{1}{2}}(d)}\tilde{\theta}_{N,q_{N}}^{1-\epsilon}(d)\to 0,\frac{1}{\sqrt{N}\Sigma_{N}^{\frac{3}{2}}(d)}M_{N}(q_{N},2)\to 0,\frac{1}{N^{2}\Sigma_{N}^{2}(d)}\sum_{p=0}^{N}|\mathcal{H}_{N}(p,q_{N})|\tilde{\theta}_{N,p}^{1-\epsilon}(d)\to 0,\vspace{-1em}$
@@
Intuitively, these additional conditions impose constraints on the expansion rate of $\mathcal{G}_{N}$, the decay in the depencence between two units, and their interactions. Consistency and asymptotic normality can be achieved when either neighborhoods in $\mathcal{G}_{N}$ grow slowly or the dependence between a unit and its remote neighbors declines sufficiently fast along the sequence $\{q_{N}\}_{N>0}$. In a geographic space, it is often reasonable to assume that two units $i$ and $j$ are no longer interfering with each other if their proximity is larger than a threshold $\bar{p}$. In this case, $\theta_{N,p}(d)=0$ if $p>\bar{p}$, and conditions 1 and 3 hold if $\Sigma_{N}(d)=O_{P}(N)$. Additionally, $\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q_{N})\right|=O\left(q_{N}^{2}\right)$, hence $M_{N}(q_{N},2)=O\left(q_{N}^{4}\right)$. Condition 2 is satisfied if $q_{N}=o\left(N^{\frac{1}{2}}\right)$, consistent with results in Ogburn et al. (2020) and Wang et al. (2025). In a network where neighborhoods expand at a polynomial rate, $\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q_{N})\right|=O\left(q_{N}^{k}\right)$, these conditions are satisfied if $\Sigma_{N}(d)=O_{P}(N)$, $q_{N}=o\left(N^{\frac{1}{k}}\right)$, and $\theta_{N,p}(d)=\frac{1}{p^{k+2}}$.
@@ 4.2 Variance Estimation
$\bar{d}_{N}=\lfloor\max\{d_{N}^{*},2d\}\rfloor,\text{ with }d_{N}^{*}=\begin{cases}\frac{1}{2}\mathcal{L}\left(\mathcal{G}_{N}\right)\text{ if }\mathcal{L}\left(\mathcal{G}_{N}\right)<2\frac{\log N}{\log\delta\left(\mathcal{G}_{N}\right)}\\ \mathcal{L}^{\frac{1}{3}}\left(\mathcal{G}_{N}\right)\text{ otherwise}\end{cases},\vspace{-1em}$
@@
[…]
@@
The second challenge is that the variance estimator may not even be conservative for $\Sigma_{N}(d)$ due to the covariance in treatment effects when interference is present, as demonstrated in both Leung (2022) and Wang et al. (2025). We solve this problem by adopting the proposal in Gao and Ding (2023). Since $\mathbf{K}_{\bar{d}_{N}}$ is symmetric, it admits an eigenvalue decomposition of the form $\mathbf{Q}_{N}\Lambda_{N}\mathbf{Q}_{N}$, where $\mathbf{Q}_{N}$ is the matrix of eigenvectors and $\Lambda_{N}$ is a diagonal matrix of eigenvalues. Gao and Ding (2023) propose replacing $\mathbf{K}_{\bar{d}_{N}}$ in Equation (9) with $\mathbf{K}_{\bar{d}_{N}}^{+}=\mathbf{Q}_{N}\max\{\Lambda_{N},0\}\mathbf{Q}_{N}$, with the maximum taken element-wise. The modified HAC variance estimator can be expressed as
$\hat{\Sigma}^{+}(d)=N\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\hat{\boldsymbol{\varepsilon}}^{\prime}_{Y}(d)\mathbf{K}_{\bar{d}_{N}}^{+}\hat{\boldsymbol{\varepsilon}}_{Y}(d)\right)\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}.$ (10)
@@
$\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)=\{j\in\mathcal{N}_{N}:d_{ij}=p\},\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=p\}\sum_{k=1}^{N}\left|K_{\bar{d}_{N},ik}\right|\sum_{l=1}^{N}\left|K_{\bar{d}_{N},jl}\right|.\vspace{-1em}$
Here, $\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)$ represents the set of unit $i$’s neighbors whose proximity to $i$ is exactly $p$. This set coincides with $\Omega_{i}(p)$ when the latter is defined as unit $i$’s $p$th-degree neighbors. $\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)$ can also be expressed as $\{(i,j,i^{\prime},j^{\prime}):j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(i,q),\,j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(j,q),\,d_{ij}=p\}$, which generalizes $\mathcal{H}_{N}(p,q)$ defined above.
@@
and $\text{and }\sum_{p=0}^{N}\mathcal{J}^{-}_{N}(p,\bar{d}_{N})\tilde{\theta}_{N,\bar{d}_{N}}(d)=O(N^{2}),\vspace{-1em}$
@@ 5 Extensions
We explore several extensions of the proposed method in this section, with technical details provided in the Supplementary Material.
@@
It is worth noting that causal identification only requires Assumption 1 to hold for potential outcomes rather than potential values of $\mathbf{L}_{i}^{(t)}$. However, the broader formulation, combined with Assumption 3, allows researchers to evaluate the assumption’s validity through placebo tests when certain variables in $\mathbf{L}_{i}^{(t)}$ are not expected to be correlated with treatment. For example, consider units with either of the two treatment histories, $(0,0)$ and $(0,1)$, in our simulated data, and suppose that $\mathbf{V}_{j}^{(t)}$ only includes variables from period $t-1$. Suppose that $\mathbf{L}_{i}^{(1)}\perp A_{j}^{(2)}\mid\mathbf{V}_{j}^{(2)}=\left(A_{j}^{(1)},\mathbf{L}_{j}^{(2)}\right)$. Then, we should expect to find a null effect of $A_{j}^{(2)}$ on the transformed outcome constructed from $\mathbf{L}_{i}^{(1)}$ using the WLS estimator. We refer to such variables in $\mathbf{L}_{i}^{(t)}$ as placebo outcomes in Algorithm 1.
@@
We consider two relaxations of Assumption 1. The first allows a unit’s treatment status in period $t$ to be influenced by the histories of nearby units, that is, treatment diffusion is permitted. We formalize the modified assumption as follows:
Assumption 7 (Sequential exchangeability with treatment diffusion)
$\Big{\{}Y_{i}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right),\mathbf{L}_{i}^{(,t+1)}\left(\mathbf{a}^{(1:T)},\mathbf{A}_{\mathcal{N}\setminus\{j\}}^{(1:T)}\right)\Big{\}}\perp A_{j}^{(t)}\mid\left(\mathbf{A}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))},\mathbf{L}_{\{j\}\cup\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}\right).\vspace{-1em}$
for any $i$, $j$, $t$, $\mathbf{a}^{(1:T)}$, and known $p$.
Here, $\mathcal{N}_{\mathcal{G}_{N}}(j,p)$ denotes the set of units whose proximity to $j$ does not exceed $p$, as defined in Section 4.1. Under Assumption 7, the variables $\mathbf{A}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ and $\mathbf{L}_{\mathcal{N}_{\mathcal{G}_{N}}(j,p)}^{(1:(t-1))}$ must be included in $\mathbf{V}_{j}^{(t)}$ and controlled for when estimating unit $j$’s propensity score at time $t$. This adjustment is straightforward when $p$ is known and fixed, and it does not affect our theoretical results. We leave the case where $p$ is unknown or grows with $N$ to future work.
The second relaxation concerns the dependence in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ across units. Assumption 1 implies that $A_{i}^{(t)}\perp A_{j}^{(t)}\mid\mathbf{V}_{\mathcal{N}}^{(t)}$ for any $i$, $j$, and $t$. However, this may not hold under assignment mechanisms like complete randomization, where treatment allocation is interdependent. In the Supplementary Material, we follow the derivation in Sävje, Aronow and Hudgens (2021) and show that our theoretical results remain valid as long as the dependence is not too strong.
@@ A Longer Period
So far, we have focused on settings where $T$ is fixed as $N$ increases. If we also allow $T$ to grow, the number of possible treatment histories expands exponentially, making it infeasible to summarize all AMRs using a saturated MSM. In such cases, researchers must impose structure on the MSM by selecting summary statistics of the treatment history and modeling their influence on the outcome. In the Supplementary Material, we show that if the MSM is saturated with respect to the chosen statistics, the resulting estimates converge to marginalized causal effects generated by those statistics. The same estimator can be used to fit the MSM, and our theoretical results still hold. Importantly, these causal quantities retain substantive meaning even if the selected statistics do not fully capture the effects of the entire treatment history.
@@
In practice, the proximity metric $d_{ij}$ may be measured with error. For example, interference might depend on commuting time or cultural similarity between two locations, yet researchers only observe straight-line distance. In social network studies, it is common to approximate the full network using self-reported ties, which often fail to capture all relevant interactions in respondents’ lives. However, such measurement error affects the definition of the estimand—that is, the particular form of spillover effect being studied—rather than its identification or estimation. In the extreme case where $d_{ij}$ is completely unavailable, one can still investigate the direct effect of a unit’s treatment history on its own outcome, accounting for interference from others in the sample. Inference becomes more challenging in such settings, as the HAC variance estimator requires researchers to specify a cutoff value $\bar{d}_{N}$. A practical approach is to assess the robustness of variance estimates across a range of $\bar{d}_{N}$ values or to use the conservative variance estimator proposed by Aronow and Samii (2017), which is built upon Young’s inequality.

### v5

@@ 4.1 Asymptotic Distribution
We restrict the dependence in $Y_{i}$ by assuming approximate neighborhood interference (ANI) introduced in Leung (2022). Let unit $i$’s $p$-neighborhood in $\mathcal{G}_{N}$ be defined as $\mathcal{N}_{\mathcal{G}_{N}}(i,p)=\{j\in\mathcal{N}:d_{ij}\leq p\}$, the set of units whose proximity to $i$ does not exceed $p$. We denote an independent copy of the treatment history outside $i$’s $p$-neighborhood as $\mathbf{A}^{(1:T)}_{i,p}=\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{\mathcal{G}_{N}}(i,p)},\tilde{\mathbf{A}}^{(1:T)}_{\mathcal{N}_{N}\setminus\mathcal{N}_{\mathcal{G}_{N}}(i,p)}\right)$. We then impose the following assumption:
@@
$\max_{i\in\mathcal{N}_{N}}E\Bigg{|}Y_{i}\left(\mathbf{A}^{(1:T)}_{\mathcal{N}_{N}}\right)-Y_{i}\left(\mathbf{A}^{(1:T)}_{i,p}\right)\Bigg{|}\leq\theta_{N,p},$
@@
This assumption implies that for any bounded Lipschitz function $L(\cdot)$, the covariance between $L\left(\mathbf{Z}_{i}(d)Y_{i}\right)$ and $L\left(\mathbf{Z}_{j}(d)Y_{j}\right)$ can be bounded by a transformation of $\theta_{N,p}$. This property, known as $\psi$-dependence (Doukhan and Louhichi, 1999; Kojevnikov, Marmer and Song, 2021), ensures that the dependence between outcomes decays as the proximity between units increases, even though it may never disappear in finite samples.
@@
$M_{N}(q,k)=\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q)\right|^{k},\vskip-10.00002pt$
@@
There exists a constant $\tilde{y}$$\infty$ such that
$\bigg{|}Y_{i}\left(\mathbf{a}^{(1:T)}_{\mathcal{N}_{N}}\right)\bigg{|}\leq\tilde{y}\vskip-10.00002pt$
@@
We discuss relaxations of this assumption in Section 5. Additionally, we assume that the propensity score can be consistently estimated at the root-$N$ rate in observational studies.
@@
We denote the normalized variance of $\hat{\boldsymbol{\beta}}(d)$, $N\times{\rm Var}\,\left[\hat{\boldsymbol{\beta}}(d)\right]$, as $\Sigma_{N}(d)$ and define $\tilde{\theta}_{N,p}(d)=\theta_{N,\lfloor p/2\rfloor}\mathbf{1}\{p>2d\}+\mathbf{1}\{p\leq 2d\}$, where $\lfloor\cdot\rfloor$ means rounding down to the nearest integer. $\tilde{\theta}_{N,p}(d)$ captures the dependence in both $\mathbf{Z}_{i}(d)$ and $Y_{i}$. We now present the main result on the large-sample properties of $\hat{\boldsymbol{\beta}}(d)$:
@@
Under Assumptions 1-6, if there exist $\epsilon>0$ and a sequence $\{q_{N}\}$ that grows with NN such that
$\frac{N^{\frac{3}{2}}}{\Sigma_{N}^{\frac{1}{2}}(d)}\tilde{\theta}_{N,q_{N}}^{1-\epsilon}(d)\to 0,\frac{1}{\sqrt{N}\Sigma_{N}^{\frac{3}{2}}(d)}M_{N}(q_{N},2)\to 0,\frac{1}{N^{2}\Sigma_{N}^{2}(d)}\sum_{p=0}^{N}|\mathcal{H}_{N}(p,q_{N})|\tilde{\theta}_{N,p}^{1-\epsilon}(d)\to 0,\vskip-10.00002pt$
@@
The proof of Theorem 1 is provided in Section B.2 of the Supplementary Material. Intuitively, these additional conditions impose constraints on the expansion rate of $\mathcal{G}_{N}$, the decay in the depencence between two units, and their interactions. Consistency and asymptotic normality can be achieved when either neighborhoods in $\mathcal{G}_{N}$ grow slowly or the dependence between a unit and its remote neighbors declines sufficiently fast along the sequence $\{q_{N}\}_{N>0}$. For example, in a geographic space, it is often reasonable to assume that two units $i$ and $j$ are no longer interfering with each other if their proximity is larger than a threshold $\bar{p}$ that does not grow with $N$. In this case, $\theta_{N,p}(d)=0$ if $p>\bar{p}$, and conditions 1 and 3 hold if $\Sigma_{N}(d)=O_{P}(1)$. Additionally, $\left|\mathcal{N}_{\mathcal{G}_{N}}(i,q_{N})\right|=O\left(q_{N}^{2}\right)$, hence $M_{N}(q_{N},2)=O\left(q_{N}^{4}\right)$. Condition 2 is satisfied if $q_{N}=o\left(N^{\frac{1}{2}}\right)$, consistent with results in Ogburn et al. (2020) and Wang et al. (2025). In a network where neighborhoods expand at an exponential rate, these conditions are satisfied if $\Sigma_{N}(d)=O_{P}(1)$, $q_{N}=O_{P}\left(\log N\right)$, and $\theta_{N,p}(d)$ decays exponentially (Leung, 2022).
@@ 4.2 Variance Estimation
$\bar{d}_{N}=\lfloor\max\{d_{N}^{*},2d\}\rfloor,\text{ with }d_{N}^{*}=\begin{cases}\frac{1}{2}\mathcal{L}\left(\mathcal{G}_{N}\right)\text{ if }\mathcal{L}\left(\mathcal{G}_{N}\right)<2\frac{\log N}{\log\delta\left(\mathcal{G}_{N}\right)}\\ \mathcal{L}^{\frac{1}{3}}\left(\mathcal{G}_{N}\right)\text{ otherwise}\end{cases},\vskip-10.00002pt$
@@
$\text{ where }\hat{\boldsymbol{\varepsilon}}_{Yi}(d)=\mathbf{Z}^{\prime}_{i}(d)Y_{i}-\sum_{k=1}^{N}\frac{\mathbf{1}\{i\in\Omega_{k}(d)\}w_{k}\mathbf{m}\left(\mathbf{A}_{k}^{(1:T)}\right)\mathbf{m}^{\prime}\left(\mathbf{A}_{k}^{(1:T)}\right)\hat{\boldsymbol{\beta}}(d)}{|\Omega_{k}(d)|}.\vskip-10.00002pt$
@@
The second challenge is that the variance estimator may not be conservative for $\Sigma_{N}(d)$ due to the covariance in treatment effects when interference is present, as demonstrated in both Leung (2022) and Wang et al. (2025). We solve this problem by adopting the proposal in Gao and Ding (2023). Since $\mathbf{K}_{\bar{d}_{N}}$ is symmetric, it admits an eigenvalue decomposition of the form $\mathbf{Q}_{N}\Lambda_{N}\mathbf{Q}_{N}$, where $\mathbf{Q}_{N}$ is the matrix of eigenvectors and $\Lambda_{N}$ is a diagonal matrix of eigenvalues. Gao and Ding (2023) propose replacing $\mathbf{K}_{\bar{d}_{N}}$ in Equation (9) with $\mathbf{K}_{\bar{d}_{N}}^{+}=\mathbf{Q}_{N}\max\{\Lambda_{N},0\}\mathbf{Q}_{N}$, with the maximum taken element-wise. The modified HAC variance estimator can be expressed as
$\hat{\Sigma}^{+}(d)=N\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}\left(\hat{\boldsymbol{\varepsilon}}^{\prime}_{Y}(d)\mathbf{K}_{\bar{d}_{N}}^{+}\hat{\boldsymbol{\varepsilon}}_{Y}(d)\right)\left(\mathbf{M}^{\prime}\mathbf{W}\mathbf{M}\right)^{-1}.\vskip-10.00002pt$ (10)
@@
$\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)=\{j\in\mathcal{N}_{N}:d_{ij}=p\},\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)=\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbf{1}\{d_{ij}=p\}\sum_{k=1}^{N}\left|K_{\bar{d}_{N},ik}\right|\sum_{l=1}^{N}\left|K_{\bar{d}_{N},jl}\right|.\vskip-10.00002pt$
Here, $\mathcal{N}_{\mathcal{G}_{N}}^{\partial}(i,p)$ represents the set of unit $i$’s neighbors whose proximity to $i$ is exactly pp. [^15] This set coincides with $\Omega_{i}(p)$ when the latter is defined as unit $i$’s $p$th-degree neighbors. $\mathcal{J}_{N}\left(p,\bar{d}_{N}\right)$ can also be expressed as $\{(i,j,i^{\prime},j^{\prime}):j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(i,q),\,j^{\prime}\in\mathcal{N}_{\mathcal{G}_{N}}(j,q),\,d_{ij}=p\}$, which generalizes $\mathcal{H}_{N}(p,q)$ defined above.
@@
$\text{and }\sum_{p=0}^{N}\mathcal{J}^{-}_{N}(p,\bar{d}_{N})\tilde{\theta}_{N,\bar{d}_{N}}(d)=O(N^{2}),\vskip-10.00002pt$
@@
The proof of Theorem 2 is provided in Section B.3 of the Supplementary Material.
@@ 5 Extensions
We briefly discuss several extensions of the proposed method in this section, with technical details provided in Section A.3 of the Supplementary Material.
@@
When certain variables in $\mathbf{L}_{i}^{(t)}$ are not expected to affect treatment, the validity of the identification assumptions can be assessed through placebo tests. For example, consider units with either of the two treatment histories, $(0,0)$ and $(0,1)$, in our simulated data, and suppose that $\mathbf{V}_{j}^{(t)}$ only includes variables from period t1t-1. Suppose that $\mathbf{L}_{i}^{(1)}\perp A_{j}^{(2)}\mid\mathbf{V}_{j}^{(2)}=\left(A_{j}^{(1)},\mathbf{L}_{j}^{(2)}\right)$. Then, we should expect to find a null effect of $A_{j}^{(2)}$ on the transformed outcome constructed from $\mathbf{L}_{i}^{(1)}$ using the WLS estimator. We refer to such variables in $\mathbf{L}_{i}^{(t)}$ as placebo outcomes in Algorithm 1.
@@
We consider two relaxations of Assumption 1. The first allows for treatment diffusion: a unit’s treatment status in period $t$ may be influenced by the histories of units whose proximity to it does not exceed $p$. When $p$ is known and fixed, researchers need only control for the histories of these nearby units when estimating the propensity scores, and our theoretical results remain unaffected. The second relaxation concerns dependence in $\mathbf{A}_{\mathcal{N}}^{(1:T)}$ across units. Assumption 1 implies that $A_{i}^{(t)}\perp A_{j}^{(t)}\mid\mathbf{V}_{\mathcal{N}}^{(t)}$ for any $i$, $j$, and $t$. However, this may not hold under assignment mechanisms like complete randomization. In the Supplementary Material, we formally state the identification assumption under treatment diffusion and, following Sävje, Aronow and Hudgens (2021), show that our theoretical results continue to hold as long as the cross-unit dependence is not too strong.
@@ A Longer Period
Our discussion so far has assumed that $T$ is fixed as $N$ increases. When $T$ also grows, the number of possible treatment histories increases exponentially, making it infeasible to summarize all AMRs with a saturated MSM. In such settings, researchers may impose structure on the MSM by using selected summary statistics of the treatment history, or focus on a shorter history, such as the most recent $s$ periods. In the Supplementary Material, we show that if the MSM is saturated with respect to the chosen statistics, the resulting estimates converge to the marginalized causal effects defined by them, even if they do not fully capture the effects of the entire treatment history. We also provide conditions under which the large-sample results continue to hold when $T$ is large.
@@
In practice, the proximity metric $d_{ij}$ may be measured with error. For example, it is common for researchers to approximate the full network using self-reported ties, which often fail to capture all relevant interactions in respondents’ lives. However, such measurement error affects the interpretation of the estimand—that is, the particular form of spillover effect being studied—rather than its identification or estimation. In the extreme case where $d_{ij}$ is completely unavailable, one can still investigate the direct effect of a unit’s treatment history on its own outcome, accounting for interference from others in the sample. Inference becomes more challenging in such settings, as the HAC variance estimator requires researchers to specify a cutoff value $\bar{d}_{N}$. A practical approach is to assess the robustness of variance estimates across a range of $\bar{d}_{N}$ values or to use the conservative variance estimator proposed by Aronow and Samii (2017).